OpenStudy (anonymous):

x-√x=2 Solve for x. The velocity of a free-falling object can be found by using the formula v=√2gd, where v is in m/s, g=9.8 m/s-squared, and d is the distance in meters. Find the distance a ball falls if its velocity is 28 m/s.

OpenStudy (diyadiya):

\[x= \sqrt{x}*\sqrt{x}\] factor out sqrt x \[\sqrt{x}(\sqrt{x} -1)=2\]\[\sqrt{x}=2 \ x=4\]Or\[\sqrt{x}-1=2 ,\ \sqrt{x}=2+1=3\]\[\sqrt{x}=3\ , x=9\] So \[x=4 \ or \ x=9\]

OpenStudy (diyadiya):

Wait i guess i did something wrong 4 is right but 9 is wrong

OpenStudy (anonymous):

Thanks.....do you know about the velocity one?

OpenStudy (diyadiya):

For the previous question its just 4 9 is wrong sorry!

OpenStudy (diyadiya):

\[v=\sqrt{2}gd\] v=28m/s g=9.8m/s^2 substitute the values and find out d

OpenStudy (anonymous):

ummm 2 time 9.8 is 4.427188724.....so do i just subtract + round?

OpenStudy (anonymous):

I'm just stuck on question 2 now.

OpenStudy (anonymous):

\[x - \sqrt{x} - 2 = 0 \]\[(\sqrt{x} - 2)(\sqrt{x} +1) = 0 \text{ , } \sqrt{x} \ge 0\] \[\sqrt{x} = 2\]\[x = 4\]

OpenStudy (anonymous):

i got that.....just the velocity question.

OpenStudy (anonymous):

v = 28 m/s = sqrt(2gd) so rearrange to find d: d = (28^2)/2g this equation comes from equating grav potential energy, mgd, with kinetic energy, (1/2)mv^2

OpenStudy (anonymous):

actually.....i got 3841.6